OpenStudy (anonymous):

78 is what percent of 82

OpenStudy (anonymous):

78=x/100 * 82

OpenStudy (anonymous):

7800=82x

OpenStudy (anonymous):

thnxs

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

"is" in math stands for = and " of "for multiplication

OpenStudy (anonymous):

type of problem

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\left(\frac{78}{82}\right)100=95.12\% \]